OpenStudy (anonymous):

15x+32y=40, x+y=10, find xy please anser this

OpenStudy (anonymous):

u mean x and y seperatly right

OpenStudy (anonymous):

no the question sent to me was exactly as I wrote.

OpenStudy (anonymous):

I guess the answer should be value of x and value of y multiplied??

OpenStudy (anonymous):

there is no mean if together

OpenStudy (anonymous):

x=280/17 , y=-110/17 xy=-30800/289

OpenStudy (anonymous):

then i must be x and y seperately, I was givin this equation by my nephew.

OpenStudy (anonymous):

is it possible to show me how you solve it. tks

OpenStudy (anonymous):

x=10-y 15(10-y)+32y=40 17y=-90 y=-90/17 x=15.29 xy=-80.968

OpenStudy (anonymous):

x+32\15y=40\32 15\32x +y=40\32 x+y=10 x+y=10 so y=32\15 Y +Y=10\32 -10 x=15\32 X+X=10\32 -10

OpenStudy (anonymous):

x=10-y 15(10-y)+32y=40 17y=-110 y=-110/17 x=10-(-110/17) x=280/17 xy=280/17*-110/17 = -30800/17

OpenStudy (anonymous):

tks all for the help, really really appreciated.

OpenStudy (texaschic101):

x + y = 10 --> x = 10 - y now sub 10 - y in for x in the other equation 15x + 32y = 40 15(10 - y) + 32y = 40 (distribute the 15 through the parenthesis) 150 - 15y + 32y = 40 (combine like terms) 150 + 17y = 40 17y = 40 - 150 17y = -110 y = -110/17 now sub -110/17 in for y x + y = 10 x - 110/17 = 10 x = 10 + 110/17 x = 170/17 + 110/17 x = 280/17 check... x + y = 10 280/17 - 110/17 = 10 170/17 = 10 10 = 10 (correct) x = 280/17 and y = -110/17 xy = 280/17 * -110/17 xy = - 30800/17 = - 1811 13/17 or - 1811.76